Vice President for Sales and Marketing Sam Totter is trying to plan for the coming year in terms of production needs to meet the sales demand. He is also trying to determine ways in which the company’s profits might be increased in the coming year. Instructions (Do all six parts): Waterways markets a simple water control and timer that it mass-produces. During last year, the company sold 701,000 units at an average selling price of 4.20 per unit. The variable expenses were $1,857,650 and the fixed expenses were $646,450. 1. What is the product’s contribution margin ratio? (Round to nearest whole percentage.) 2. What is the company’s break-even point in units and in dollars for this product? 3. What is the margin of safety, both in dollars and as a ratio? (Round to nearest whole percentage.) 4. If management wanted to increase its net operating income from this product by 10%, how many additional units would have to be sold to reach this income level? 5. If sales increase by 51,000 units and the cost behaviors do not change, how much will net operating income increase on this product? 6. Waterways’ management believes that increased advertising would increase unit sales by 10%. If management wants to increase its operating income by $50,000, how much could the company spend on additional advertising to reach its goal?
Answer #1

1)

Sales $ 2,944,200.00
Variable Expense $ 1,857,650.00
Contribution Margin $ 1,086,550.00
contribution margin ratio 37%

2) Break Even Point in Dollars = 646,450/37%

=1,747,162

Break Even Point in units = 1,747,162/4.2

=415,991 Units

3)margin of safety in Dollars = 2,944,200 - 1,747,1622

=1,197,038

margin of safety Ratio =1,197,038/2,944,200

=41%

4)Additional income = 44,010

Additional Sales Required in Dolars = 44,010/37% = 118,946

In Units = 118,946/4.2

=28,320 Units

5)51,000 x 37% x 4.2 = 79,254

6)Target Operating Income = 490,100

New Sales = 3,238,620

Contribution Margin = 3,238,620 x 37% = 1,198,289.40

Fixed cost + Target income = 646,450 + 490,100

=1,136,550

additional advertising = 1,198,289.40 - 1,136,550

=61,739.40